3d²y/dx²+dy/dx-14y=0y(0)=1y(0)=-1


1
Expert's answer
2022-03-15T15:09:58-0400
"3\frac{d²y}{dx²}+\frac{dy}{dx}-14y=0"

The characteristic equation is as follows

"3k^2+k-14=0"

Roots

"k_1=-7/3,\quad k_2=2"

General solution of DE

"y(x)=C_1e^{-7\/3x}+C_2e^{2x}"

Using initial condition, we obtain


"y(0)=C_1+C_2=1\\
y'(0)=\frac{-7}{3}C_1+2C_2=-1"

Roots:

"C_1=9/13,\quad C_2=4/13"

Finally, the particular solution of DE


"y(x)=9\/13e^{-7\/3x}+4\/13e^{2x}"